OpenStudy (anonymous):

Jennie purchased 5 of the 200 tickets at the raffle. There will be two drawings for a prize. The probability that Jennie will win both of the prizes is 1 out of ____. 1,592 1,600 1,990

OpenStudy (auctoratrox):

(1/40)(1/40) = 1/1600

OpenStudy (anonymous):

P (winning ticket) = 5/ 200 = 1/40 P ( 2 winning tickets) = ( 1/40 ) ^2 = ...

OpenStudy (anonymous):

Probability= 5/200 x 4/199 =1/1990 1st drawing there are 5 tickets which are Jennie's and a total of 200 tickets. So after drawing the 1st ticket, 4 tickets are Jennie's and left with a total of 199 tickets to draw, hence the 4/199.